# SelfAI: Building a Self-Training AI System with LLM Agents

### v1

@@ 2.3 Computer Vision
2.4 Medical Image Analysis
In medical image analysis, nnU-Net [72] is a landmark framework known for its strong generalization and automated design. Despite its adaptive network structure, preprocessing, and training strategies for various segmentation tasks, the exploration of novel architectures persists. To address this, nnU-Net-Revisited [73] establishes a comprehensive benchmark including 19 mainstream models (CNN-based [74, 75, 76], Transformer-based [77, 78], and Mamba-based [79, 80]), offering a solid basis for fair evaluation. Results are shown in Figs. 2-3 and Supplementary Tables 1213. On the BTCV dataset [81], GPT4-o3-mini achieves the highest Score (0.6875) and Rank 1, demonstrating strong optimization efficacy, while Qwen2.5-7B secures second place, outperforming all larger variants in its family. In nnU-Net hyperparameter tuning for the BraTS dataset [82], Qwen2.5-14B attains the top ranking (Score 0.4333), followed closely by GPT4-o3-mini, with Qwen2.5-32B and DeepSeek-R1-70B tying for third. Notably, the 7B and 72B models outperformed similar models of medium size (14B–32B), indicating that larger model size does not guarantee better optimization.
2.5 Imbalanced Node Classification
Graph neural networks (GNNs), as powerful tools for processing relational data, are playing an increasingly critical role in numerous cutting-edge scientific fields. In biomedicine, GNNs have become a core component of protein structure prediction models such as AlphaFold3 [39], enabling precise modeling of spatial interactions between amino acid residues and ushering in a new era in structural biology. In drug discovery, they are widely used for molecular property prediction [39, 83, 84], compound-target interaction identification [85, 86], and novel drug generation [87]. Similarly, GNNs demonstrate exceptional performance in engineering domains, including traffic prediction [88, 89], cybersecurity, and chip design [90, 91]. Given this broad utility and the sensitivity of GNN performance to hyperparameter settings, we evaluate GraphSAGE on the imbalanced Cora benchmark [92] under the setup of [93].
Figs. 2 and 3 represent that the GPT4-o3 series delivered near-optimal results, achieving the fastest discovery of optimal solutions and efficient stopping decision while maintaining high search diversity, underscoring its robust optimization capability in scientific computing scenarios. In contrast, Qwen2.5 models, (except 72b), easily delivered consistent but middling results across scales, while larger models generally exhibit late convergence and higher unnecessary exploration, reflecting potential inherent limitations in understanding the structural properties and optimization requirements of specialized domains such as protein interaction networks and molecular graph structures (see Supplementary Table 14 for more results). These results indicate that SelfAI can effectively address challenges in GNN hyperparameter optimization. The framework’s strong performance on graph learning tasks suggests its potential as a general-purpose optimizer for scientific domains, including protein engineering and drug development.
2.6 Drug Discovery
In drug discovery, accurate bioactivity prediction is vital for early-stage virtual screening and compound prioritization, enabling shorter development cycles and lower preclinical costs. Following the evaluation practices and modeling standards summarized by Korotcov et al. [94], this study employs the Chagas EP20 dataset [95], which measures the activity of compounds against Trypanosoma cruzi, a neglected but medically significant target in tropical disease research. The core challenge lies in learning an effective mapping from molecular representations, such as SMILES sequences [96] or graph-based encodings [97] to experimentally measured bioactivity [98]. We adopt SelfAI to adaptively refine performance, assess search trajectories, and determine principled stopping points for this task.
Experimental results (see Supplementary Table 15) reveal distinct performance patterns among solvers in hyperparameter optimization. Qwen2.5-7B and 14B secure first and second place, respectively, both achieving high Scores with minimal search diversity. GPT-4o ranks third, while DeepSeek-R1-7B and 70B tie for fourth. A key finding is the clear performance drop at larger scales; Qwen2.5-32B, Qwen2.5-72B, and DeepSeek-R1-32B rank in the bottom half, demonstrating that increased parameters do not guarantee better optimization. In contrast, traditional optimization methods (such as GS and BS) ranked poorly in this task, underscoring the clear advantage of large language model-based intelligent optimizers in bioactivity prediction. These results collectively demonstrate that in high-dimensional, noisy biochemical scenarios, language model-driven optimizers offer efficient convergence and a superior balance between exploration and exploitation.
Finally, the ranking heatmap (Fig. 7) and average performance analysis (Supplementary Table 3) reveal several consistent trends. GPT-4o-mini emerges as the most consistently effective solver, strong inductive reasoning, rapid adaptation to emerging evidence, and stable progression across diverse scientific domains. Mid-sized models such as Qwen2.5-7B and Qwen2.5-14B display competitive and robust performance relative to their scale, often exhibiting disciplined evidence-based refinement during exploration. In contrast, larger models (Qwen2.5-72B and DeepSeek-R1-70B) exhibit higher variance: they perform well on specific tasks but frequently show delayed breakthroughs, unstable trajectories, and inconsistent cumulative progress. These patterns suggest that larger models tend to spend more time exploring alternative possibilities rather than leveraging early evidence, resulting in delayed commitment and reduced adaptability during the search process.
@@ 3 Discussion
The emergence of the autonomous scientific discovery system (ASDS) represents a paradigm shift in scientific research, with the potential to accelerate discovery by delegating structured experimental processes to artificial intelligence. Yet most existing agentic ML frameworks and benchmark-driven ASDS, such as MLE-bench [99], emphasize an agent’s ability to generate and execute candidate solutions, offering limited support for adaptive trajectory management, such as reasoning about when exploration should continue, how search strategies should evolve, or why certain regions of the space warrant further investigation. SelfAI addresses this gap by integrating structured intent interpretation, multi-agent reasoning, and principled stopping criteria into a general reasoning framework. Across diverse domains, this design enables the system to allocate computational effort where it is most informative, rather than exhaustively enumerating the search space.
Under this goal, SelfAI is meticulously designed as a general-purpose platform capable of performing high-level reasoning across diverse tasks, thereby accelerating the experimental search process. Evaluations across a comprehensive benchmark (12 tasks across 6 different domains) uncover that SelfAI strategically allocates limited computational resources to the most promising performance spaces or research directions and terminates unproductive trajectories, exhibiting human-like cognitive flexibility in dynamically adjusting the search trajectory. This optimal stopping behavior effectively alleviates the most persistent inefficiencies in current ASDS.
The newly proposed Score and $\text{AUP}_{D}$ metrics are crucial for quantifying this advantage, revealing SelfAI’s remarkable balance between rapidly finding discovery efficiency (Score) and maintaining healthy exploration diversity ($\text{AUP}_{D}$). These metrics further illuminate the limitations of traditional methods like grid search and Bayesian optimization, which often lack the adaptive reasoning required to escape local optima or terminate unfruitful trials. Our findings also suggest that in scientific domains, effective scientific discovery demands specialized reasoning capabilities beyond mere model scale. Although large-scale models like GPT-4o and DeepSeek-R1-70B occasionally excel in broad exploration, their performance is frequently surpassed by more compact, purpose-driven models such as Qwen2.5-7B and DeepSeek-R1-7B, which demonstrate superior stopping efficiency and search consistency. This indicates that for navigating complex non-convex optimization environments, the structured reasoning and hypothesis generation implemented in SelfAI’s cognitive agent may play a more decisive role than model scale alone. The framework demonstrates that purpose-driven architectural design can effectively harness the capabilities of moderately-sized models to deliver robust scientific discovery performance.
The implications of this framework extend far beyond accelerated experimental search and improved benchmark performance. First, the architecture of SelfAI embodies a practical blueprint for human-AI symbiosis in science. The User Agent acts as a seamless translator, allowing human researchers to interact with the system at the level of scientific intent and high-level questions, rather than through low-level code or configuration files. This design formally integrates human expertise as a guiding input, creating a collaborative loop where the AI manages the scale and complexity of the search while the human provides strategic direction and domain insight. This addresses a critical shortcoming of "black-box" automation by fostering reproducibility and trust, as every action of the AI is traceable to a human-defined objective. Furthermore, this closed-loop process allows researchers to shift from using AI for scientific research to exploring how to conduct scientific research optimally and to investigate efficient search strategies and problem-related endpoints under controlled conditions.
SelfAI has the potential to evolve from a tool into a collaborative research partner, enhancing the cognitive and experimental capabilities of human scientists. This evolution would be driven by several key advancements. For instance, the User Agent could be enhanced to incorporate autonomous code generation, thereby broadening the range of experimental configurations without requiring explicit user pre-definition. Furthermore, while the current iteration of SelfAI relies on the static knowledge embedded within its underlying LLM (Large Language Model), future versions could leverage advanced memory management and RAG approaches to alleviate context length limitations and dynamically integrate a richer, more up-to-date body of domain-specific knowledge. Collectively, these envisioned capabilities underscore SelfAI’s foundational role in shaping a more adaptive, collaborative, and intelligent paradigm for autonomous scientific discovery.
4 The Design of SelfAI for Scientific Discovery
SelfAI is a general scientific discovery system designed to automate and accelerate the end-to-end scientific discovery process. Its core function is to transform users’ high-level research intentions (hyperparameter optimization, novel algorithm design, model architecture validation, or ablation studies) into structured, executable workflows. The system operates in an integrated loop, comprising three core agents: a user agent, a cognitive agent, and an experiment manager, which together constitute the overall experimental workflow in Fig. 1.
4.1 User Agent
The User Agent, as the user interface of SelfAI, is designed to be interactive, user-controllable, and adaptable to evolving research needs. It interprets high-level scientific intent and translates natural-language objectives into structured, machine-readable experimental configurations. Through iterative clarification, it helps researchers specify objectives, constraints, and search spaces, thereby establishing the experimental environment in which the Cognitive Agent operates. For example, requests such as “achieve state-of-the-art accuracy on CIFAR-10 using a CNN” or “identify the most influential method for image classification” are reformulated into precise experimental specifications (see Supplementary Material A.1). Unlike static intent-generation prompts, the User Agent supports continuous intervention: researchers may pause execution, inspect reasoning traces, adjust constraints, or redesign experimental protocols at any stage without restarting the workflow.
4.2 Cognitive Agent
SelfAI’s cognitive agent is central to accelerating scientific discovery. After receiving structured experimental configurations from the user agent and accumulated experimental history from the experiment manager, the agent synthesizes all available information, analyzes experimental trajectories, evaluates observed performance trends, and identifies promising regions in the search space. This achieves a delicate balance between developing these high-potential regions and exploring areas of greater uncertainty, ultimately deriving the next optimal search strategy. A significant feature of our cognitive agent is its introduction of principled stopping criteria, addressing a key limitation of traditional AI systems, which often continue exploring under conditions of diminishing returns. By conducting evidence-based evaluations of research progress, the agent can determine when further experiments on a particular hypothesis are unlikely to yield significant new insights, thereby reallocating resources to more promising research directions (see Supplementary Material A.2 for more details). This capability improves resource utilization efficiency in the automated discovery process.
4.3 Experiment Manager
The Experiment Manager executes the experimental plans proposed by the Cognitive Agent, each of which has already been structured as executable training jobs, specifying model architectures, datasets, and hyperparameters. For every job, the manager initializes these specified configurations, launches the training and evaluation pipeline, and records all associated outputs. It tracks runtime status, handles execution failures, and captures performance metrics as well as relevant logs and metadata. These records constitute a continuously updated structured knowledge base, providing a reliable foundation for the cognitive agent to analyze search progress, identify promising regions of the search space, and decide when to stop exploration. Comprehensive execution procedures and implementation details are provided in Supplementary Material A.3.
Beyond executing trials, the Experiment Manager also manages the computational environment in which scientific experiments are carried out. Through the Experiment Manager, resources are allocated reliably, execution workloads are scheduled efficiently, and checkpointing and recovery mechanisms are coordinated seamlessly. Anomalies or failed runs are surfaced to the Cognitive Agent as feedback for complex adaptive reasoning. As a result, the overall research process is conducted within an efficient, reliable, and reasoning-driven environment.
@@ 4.3 Experiment Manager
Functions Ours Code LLaMA [53] MLGym [48] AI Co-Scientist [45] AIRA [24] MLAgentBench [49] Optuna [61]
@@
4.4 Comparison of AI research systems
We contextualize SelfAI’s capabilities within the current landscape of AI-assisted research tools. Table 1 presents a systematic comparison against representative frameworks across multiple dimensions of scientific functionality. Existing ASDSs often specialize in code- or idea-level hypothesis generation under fully autonomous pipelines that primarily optimize medal rates within 24 hours. In contrast, SelfAI integrates intent interpretation, iterative scientific reasoning, and adaptive trajectory optimization into a unified discovery workflow. These capabilities allow the system to refine strategic planning based on accumulated experimental evidence, rather than executing predetermined and static generalization procedures. Overall, these advantages position SelfAI as a general-purpose framework capable of supporting efficient, reliable, and reasoning-driven scientific workflows that extend the scope of existing agent or hyperparameter optimization benchmarks.
5 Data availability
All datasets used in this study are available for download at https://github.com/XiaoXiao-Woo/SelfAI.
6 Code availability
The SelfAI repository is available as Supplementary Software. Updated versions can be found at https://github.com/XiaoXiao-Woo/SelfAI.
Appendix A LLM Agents in SelfAI System for Scientific Discovery
A.1 User Agent: Idea Interaction and Experiment Configuration
In autonomous research systems, the User Agent is typically powered by a general-purpose LLM and acts as the primary interface between human researchers and the system. These systems build an “idea-to-experiment” autonomous research process where user access to detailed experimental workflows is limited. Such approaches can streamline research execution but may also restrict researchers’ flexibility in iteratively shaping and adapting experiments. By contrast, our User Agent unifies idea generation and experiment configuration into a continuous, interactive process, empowering researchers to refine initial concepts, explore alternative approaches, and adjust experimental parameters at any stage.
During the idea interaction process, the User Agent understands the task background, dissects method details, and defines precise experimental objectives across diverse scientific domains. At this stage, the User Agent can help users pinpoint innovative research directions and generate field-relevant ideas. Once an idea emerges, the configuration interaction phase leverages prompt templates to clarify user objectives and formulate strategic experimental plans, including initial hypotheses and search configurations. This process converts high-level manual queries into standardized and reproducible experimental configurations, as illustrated in the following examples. We insert the specified template into the prompt and trigger the prompt to fill out the specified configuration file.
Consequently, the User Agent can transform ideas into system and user information. The system information provides the Cognitive Agent with essential content, including role-playing instructions and task-specific knowledge (such as task benchmarking, method evaluation, comparative analysis, ablation studies, and experimental designs). This information equips the Cognitive Agent to interpret high-level research objectives accurately and to execute relevant reasoning and planning steps effectively.
In parallel, the user information focuses on the practical aspects of experimental planning by supplying detailed specifications and keywords, such as search space definitions, the number of trials, optimization criteria, and resource constraints, which are critical for configuring and running experiments. Together, these two types of information enable seamless coordination between conceptual ideation and concrete execution, supporting a workflow in which researchers can iteratively refine their experiments and monitor progress in real time.
A.2 Cognitive Agent
Figure 5: Illustration of the Cognitive Agent. The overall reasoning process involves several key steps: Hypothesis Generation (analysis of the current task and completed trials), Stopping Judgment, and Strategic Planning. Strategic Planning develops experimental schemes based on the analyzed hypotheses.
In our SelfAI framework, the Cognitive Agent continuously ingests and adapts domain knowledge to tackle complex reasoning problems during experiments, much like a human researcher. At each stage, the agent first conducts a comprehensive analysis of the task and completed experiments to clarify the purpose and key information of the task and generates preliminary hypotheses. Then, it evaluates whether these hypotheses are worth trying through stop judgment to inform strategic planning. It then generates specific schemes to probe unexplored regions of the solution space, creating a non-Markovian chain of evolving plans grounded in prior states.
A.2.1 Reasoning for Scientific Discovery
LLM agents exhibit remarkable reasoning capabilities, making them well-suited for scientific discovery. As illustrated in Fig. 5, we present an advanced reasoning framework for generating optimal stopping solutions, an approach that can be extended to various scientific discovery tasks.
A. Trajectory Analysis and Hypothesis Generation
Specifically, the agent begins with Task 1: Analyze the Current Task, where it meticulously interprets the prompt to establish the task’s core objectives, constraints, and hyperparameters. This initial analysis is used to perform a structured trajectory evaluation, creating a coherent chain of thought that guides all subsequent strategy and exploration. This initial analysis is used to perform a structured trajectory evaluation, creating a coherent chain of thought that guides all subsequent reasoning and planning strategies.
B. Best Stopping Judgement As described in Sect. LABEL:sec:self-evaluation, we introduce the optimal stopping criterion to guide the prompt design that balances exploration and exploitation. During Best Stopping Judgement, the cognitive agent first determines whether the current performance metrics surpass those of the initial configuration and conducts an optimal stopping judgment. Building on the prior analysis of completed trials, the agent systematically evaluates stopping conditions to avoid testing all configurations.
C. Strategic Planning Following the reasoning and stopping judgment phases shown in Fig. 5, the agent reviews all executed trials to synthesize key observations on performance trends, notable configurations, and areas of consistent improvement. Next, it translates its abstract hypotheses into a concrete experimental plan. This strategic process balances two competing objectives: exploiting promising configurations and exploring uncertain regions. Exploitation focuses on refining or building upon previously successful configurations, whereas exploration targets under-examined areas that show significant potential for performance gains. Each proposed trial is explicitly justified by insights from the preceding reasoning phase, ensuring the search is deliberate and well-founded. Finally, each recommendation is formalized as an intended trial and delegated to the Experiment Manager for execution.
As trials advance and incremental improvements diminish, the Experiment Manager maintains a dynamic record of completed actions and continuously refines the remaining search space. Strategic planning directs the optimization trajectory toward convergence, and the system employs an effective stopping strategy to terminate exploration in a rational and efficient manner.
Notably, the LLM-based agent demonstrates human-like exploratory behavior: it initially focuses on refining hypotheses around high-performing regions and gradually extends attention to areas of greater uncertainty. This pattern reflects a synergy of intelligence-driven reasoning and adaptive planning.
A.3 Experiment Manager

### v2

@@ 3 Discussion
Autonomous scientific discovery systems (ASDS) have been widely proposed as a means to accelerate research by delegating large-scale experimentation to artificial intelligence. However, most existing agent frameworks and ASDS benchmarks primarily emphasize candidate generation and execution, offering limited support for managing long-horizon discovery processes. In particular, they provide little explicit reasoning about how search strategies should evolve over time, when exploration should be terminated, or how experimental effort should be redistributed as evidence accumulates.
This work addresses these limitations by reframing scientific discovery as a trajectory-level decision-making problem. By integrating structured intent interpretation, trajectory-aware reasoning, and adaptive stopping within a unified framework, SelfAI enables explicit reasoning over accumulated experimental evidence. Across diverse benchmarks, SelfAI consistently reallocates computational resources toward informative regions of the search space and terminates exploration under diminishing returns, reshaping discovery dynamics beyond what is achievable through endpoint-focused optimization alone.
The newly proposed trajectory-aware evaluation metrics are crucial for quantifying this advantage. Metrics such as Score and $\text{AUP}_{D}$ reveal substantial differences in discovery efficiency, redundancy, and stopping behavior that are obscured by conventional endpoint-based evaluation. While many solvers ultimately converge to similar final performance, trajectory-level metrics expose how quickly high-quality configurations are identified and how effectively redundant experimentation is avoided. These measurements highlight structural limitations of classical optimization methods, including grid search and Bayesian optimization, which lack adaptive reasoning mechanisms for long-horizon exploration and principled termination.
SelfAI’s trajectory-aware reasoning enables long-term scientific exploration by integrating experimental history. While our benchmark results show that LLM-based agents generally exhibit broad exploration behavior and consistent final performance, they can also fall into the trap of excessively long hypothesis exploration times, model execution delays, and greater performance fluctuations. Although some small- to medium-sized models, guided by structured reasoning and adaptive stopping, demonstrate more stable trajectory organization and earlier convergence, optimal stopping decisions can still fail in highly complex environments. Premature stopping can still lead to the inability to identify the global optimum, while over-exploration can occur under extreme uncertainty (see Supplementary Sect. B.6). To enhance the long-term scientific exploration capabilities of LLM agents, dynamic knowledge integration and contextual learning capabilities can be integrated to overcome the inherent knowledge bottlenecks of LLMs.
Beyond benchmark performance, the proposed framework has broader implications for human-AI collaboration in scientific research. By separating high-level scientific intent from low-level experimental execution, SelfAI provides a structured interface through which human expertise can guide large-scale automated exploration. This design promotes transparency and reproducibility by ensuring that exploration policies remain anchored to explicit research objectives. More broadly, it enables a shift from using AI solely as an optimization tool toward studying scientific discovery itself as an object of inquiry—allowing researchers to analyze, compare, and refine discovery strategies under controlled conditions.
Finally, LLM-based agents enable explicit modeling of long-horizon scientific discovery, shifting discovery systems from task-specific performance optimization toward continuous discovery processes for broader scientific domains. Considering a wide range of scientific fields, including physical and materials sciences, there is broad interest but limited expertise in general-purpose LLM agents. SelfAI can interleave LLM agents from different application domains to support the design of entirely new systems. Moreover, this framework can be integrated with automated toolchains, such as code generation and more powerful inference models, to enable smarter and more adaptive discovery workflows. Overall, this work opens new opportunities for scientific discovery by establishing a pathway toward continuous, long-horizon exploration.
4 Materials and Methods
4.1 Scientific discovery benchmark
Existing machine learning benchmarks, such as MLE-bench [chan2024mle], are primarily designed to accelerate the engineering of learning systems by optimizing performance under constrained computational budgets. In contrast, we introduce a scientific discovery benchmark specifically tailored to long-horizon discovery, formulated as a trajectory-driven decision-making process. Rather than emphasizing final performance, the benchmark requires solvers to engage in scientific reasoning by analyzing experimental outcomes, interpreting accumulated observations, and adaptively revising search strategies to uncover functional relationships, structural patterns, or optimal experimental configurations.
The benchmark comprises 12 tasks spanning six scientific categories, including scientific computing, machine learning, computer vision, medical image analysis, and drug discovery (Table 1). All benchmark data are collected from real experimental runs, with several tasks corresponding to configurations reported in published literature. The tasks cover both discrete and continuous search spaces and range from low- to high-dimensional settings. To simulate realistic user-system interaction, each task provides background information to the LLM to support intent interpretation. Most datasets are generated through systematic grid-based exploration, while LCBench and the Chagas EP20 drug discovery dataset are constructed from Bayesian optimization trajectories. LCBench is a widely used AutoML benchmark [zimmer2021auto]. All tasks are organized as YAML specifications containing metadata, task definitions, concise problem descriptions, search spaces, and trial records. The benchmark will be publicly released at https://github.com/XiaoXiao-Woo/SelfAI.
4.2 Compared methods and evaluation setup
We evaluate SelfAI across all 12 benchmark tasks and compare it with classical optimization methods, including grid search and the Tree Parzen Estimator (TPE) optimizer [optuna], referred to as BS. To assess robustness across model families, we instantiate SelfAI with different language-model backbones, including OpenAI-o3 [openai2023gpt], Llama3.3 [llama], Qwen2.5 [bai2023qwen], and DeepSeek-R1 [guo2025deepseek]. All methods are evaluated under identical computational settings. Reasoning traces are retained for analysis and reproducibility. For all trials, the random seed is fixed, and the temperature is set to zero. The detailed settings of the benchmark are available in the Supplementary Materials.
Category Method Task Dim Count Ref.
Scientific Computing Tensor Network Image Completion 3 64 [tw]
Machine Learning Random Forest Pricing Prediction 5 162 [boston]
LSTM Sentiment Analysis 2 20 [sentiment_analysis]
GraphSAGE Node Classification 22 25 [yan2023unreal]
Computer Vision SIREN Image Denoising 2 25 [sitzmann2020implicit]
SIREN Image Segmentation 2 25 [sitzmann2020implicit]
ResNet Image Classification 4 9 [resnet]
MAE Image Classification 2 20 [mae]
FashionMnist-NN Image Classification 5 2000 [zimmer2021auto]
Medical Image Analysis nnUnet BraTS [brats] 3 18 [nnunet]
nnUnet-revisited BTCV [btcv] 5 19 [nnunet_revisit]
Drug Discovery DNN Bioactivity Prediction 4 30 [korotcov2017comparison]
4.3 SelfAI system
The SelfAI system operates in an integrated loop, comprising three core agents: User Agent, Cognitive Agent, and Experiment manager, which together constitute the overall experimental pipeline in Fig. 1. It consists of three components: a User Agent for intent formalization, a Cognitive Agent for iterative decision-making, and an Experiment Manager for experiment orchestration and management. Before illustrating the details of SelfAI, we first situate SelfAI within the landscape of existing AI-assisted research frameworks by examining their functional coverage (Table 2).
@@ 4.3 SelfAI system
Functions Ours Code LLaMA [kochnev2025optuna] MLGym [MLgym] AI Co-Scientist [ai_co_scientist] AIRA [aira] MLAgentBench [mlagentbench] Optuna [optuna]
@@
4.3.1 User agent
As illustrated in Fig. 1b, User Agent serves as the user interface through which high-level research intent is translated into structured, machine-readable experimental specifications. Its primary function is to formalize natural-language objectives provided by users into a standardized configuration that defines the experimental task for the SelfAI system. For example, user requests such as “Design high-performance deep learning for image classification task on the ImageNet dataset” or “Identify the most influential method for image classification” are reformulated into precise experimental specifications. Specifically, User Agent maps user-provided descriptions of research goals, constraints, and preferences into a task-specific YAML configuration. This configuration explicitly specifies basic ideas, optimization objectives, the experimental search space, trial budget, and historical trial data (as shown in Supplementary Sect. A.1). The resulting YAML file serves as the sole interface between user intent and downstream system components, including Cognitive Agent and Experiment Manager. Consequently, User Agent encodes the user’s research intent into a fixed configuration schema, which is implemented via a predefined prompt template. All interactions are restricted to the generation and refinement of the structured configuration, and users do not directly manipulate experimental parameters or outcomes once the configuration is finalized.
4.3.2 Cognitive agent
Cognitive Agent performs iterative decision-making within the SelfAI framework by operating on structured experimental inputs and conducting trajectory-level analysis across multiple iterations. To support long-horizon optimization, Cognitive Agent operates in three functional stages: hypothesis generation, strategic planning, and stopping judgment. Hypothesis generation identifies promising regions of the search space based on accumulated experimental outcomes. Strategic planning translates these hypotheses into concrete experimental proposals that balance refinement of high-performing configurations with exploration of under-sampled or uncertain regions. Stopping judgment evaluates accumulated experimental evidence to determine whether continued experimentation is warranted under the current exploration strategy. At each iteration, Cognitive Agent receives a task-specific experimental configuration from User Agent, along with accumulated experimental histories and performance metrics from Experiment Manager. These inputs are provided in a structured format and define the current state of the search process. Using the accumulated evidence, Cognitive Agent evaluates experimental trajectories, assesses observed performance trends, and estimates coverage of the explored search space. Based on this analysis, it generates candidate experimental configurations for subsequent exploration. In addition to proposing new experiments, the Cognitive Agent produces a stopping decision based on trajectory-level evaluation of completed trials, including comparison with the initial configuration, assessment of diminishing returns in recent performance improvements, and estimation of the potential value of remaining unexplored regions. The outputs of Cognitive Agent consist of candidate experimental configurations for the next iteration and a stopping signal indicating whether the current exploration process should be terminated. Detailed implementation, including the prompt templates used for hypothesis generation, strategic planning, and stopping judgment, is provided in the Supplementary Materials A.2.
4.3.3 Experiment manager